B is 50% faster than A. If A starts at 9 A.M. and B starts at 10 A.M. A travels at a speed of 50 km/hr. If A and B are 300 kms apart, The time when they meet when they travel in opposite direction is ?
Answer by lwsshak3(11628)

B is 50% faster than A. If A starts at 9 A.M. and B starts at 10 A.M. A travels at a speed of 50 km/hr. If A and B are 300 kms apart, The time when they meet when they travel in opposite direction is ?
***
By 10 A.M., A had been traveling for 1 hr, covering a distance of 50 Km at 50 km per hr.
Speed of B=50+50%*50=50+25=75 km/hr
let x=travel time (hours) it would take for A & B to meet after 10 A.M.
(x+1) = total travel time (hours) of A
distance=speed*travel time
50(x+1)+75x=300
50x+50+75x=300
125x=250
x=2 hrs
Ans:
The time when they meet when they travel in opposite directions is=10+2=12 noon